Oak Farms is an unlevered firm with 3055 shares outstanding and an EBIT of 650. Corporate earnings are taxed at a rate of 32%. Calculate EPS for Oak Farms. $ Note: Your answer should be in dollars and cents. For example, $0.99. Suppose that Oak Farms makes a decision to partition (split) its assets into debt and equity. The firm issues $2150 of debt at a cost of 8.85%, and uses these funds to reduce the amount of equity on its books. The partition does not change the EBIT or the tax rate, but does reduce the number of shares outstanding to 2305. Compute Oak Farms EPS after the partition. $ Note: Your answer should be in dollars and cents. You Must Get Both Parts Correct to Receive Credit
Answer #1

The solution to the given question is provided below for your reference

1.
=650*(1-32%)/3055=0.144680851

2.
=(650-2150*8.85%)*(1-32%)/2305=0.135623861